GreenHab Report – Mar 15th

Crew 206 Green Hab Report – 15-MAR-2019

GreenHab Officer: Norbert Pouzin

Environmental control: Heating and Cooling w/ ambient air (10 hrs)

Shade Cloth (40%)

Average temperatures:
Low: 17.5°C
High: 30.5°C

Hours of supplemental light: 0 Hours

Daily water usage for crops: 6.4 gallons

Water in Blue Tank: 198.0 gallons

Crops watered at 0810 hours (3.2 gallons) and 1930 hours (3.2 gallons):

TIME NOTED TEMPERATURE(° C ) / HUMIDITY(%)

1) 08.10 AM 26.1° C / 23 %
2) 12.00 AM 26.2° C / 20 %
3) 15.00 PM 27.8° C / 20 %
4) 19.30 PM 23.5° C / 25 %

Changes to crops: Sunflowers were removed.

Narrative:
Door was opened today from 8h to 18h. I finished my experiment today and cleaned the greenhab. Mission is over, everything went pretty well in this peaceful and beautiful place. I will miss these moments in the greenhab. My experiment showed me that music can help monitor plants (here radishes), and it gives me new ideas for my return on Earth !

Harvest: 14 g of basil, 5 g of sage, 1 g of origano, 10 g of parsley.

Support/supplies needed: None.

Journalist Report – Mar 15th

Sol 19

Authors: Benjamin Auzou, Journalist, & Norbert Pouzin, Greenhab Officer

The end of an adventure but not the end of the story

On the 24th February, we closed the airlock of the station and started our mission. Three weeks after, we opened the airlock and ended our rotation in the MDRS. These 20 Sols on Mars were intense and unprecedented for us. At 18:01 we opened the main airlock door and ran on the Earth’s ground without a suit for the first time since the beginning of the simulation. Even if we were happy in the Hab, it’s good to be back on Earth and breathe the fresh air. Feeling the atmosphere, or walking on the ground, things appearing to be basic on
everyday life was something unique tonight.

Mars is an ambitious goal for humanity, but Earth is our cradle and feeling it again was more than a simple pleasure. We realized that humanity needs Earth to survive and that we have to take care of our Pale Blue Dot. These three weeks here taught us that our lives don’t need as much comfort as they have. A shower everyday, over-connections, and over-consumption don’t give us as much pleasure as we consume and destroy our planet to have it.

Living here on Mars showed us what was really essential to live and what is important in human relationships. A bit of cooking, nice friendships, great experiments and pétanque seem enough. However, Earth calls us back ; we will soon come back to our daily routines in
France, but we will never forget what happened in the last weeks.

We met on August 2017 in Toulouse, Jérémy was in second year of ISAE-SUPAERO and the rest of the crew arrived for its first year. Then we started working together around the projects of the student association Club MARS. Our group that became crew 206 gets stronger days after days, as we were working together with a same goal, this mission. After more than a year to prepare it, facing many issues, joy, difficulties, we have achieved our mission. We are all very happy for what has been done, for the moments all together, and the science we have made. If the mission is over, what we have achieved is just starting.

Aurélien, Norbert and I were the last marswalkers of our mission. We walked through Candor Chasma, enjoying our last steps and our last views of Mars. We lived a wonderful adventure here and participated actively to take the scientific challenges of the exploration of Mars and we warmly thank the Mars Society, Robert Zubrin, Shannon Rupert, Atila Mezsaros and all the team of CapCom for making this possible.

We spent three extraordinary weeks that we will never forget and that taught us a lot of things. We are more than a crew; we are six friends with strong links that are even bigger than this mission.

Operations Report – Mar 15th

Crew 206 Operations Report 15 – MAR – 2019

SOL: 19

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and
alternator need replacement on generator. A contractor has
been arranged to make these repairs within a few days.

Generator (hours run): ??? hours

Solar— SOC % – Turned off at ??%, Still off at 90% (1930)

Diesel Reading – 55%
Propane Reading – 70%

Ethanol Free Gasoline – 2 gallons

Water (Auxillary tank) – Not in use
Water (Static tank) – 300 gallons
Water differential (static tank) – 12 gallons

Gallons transferred: Not Applicable

Water in GreenHab – 198 gallons

Static to Loft Pump used – Yes
Water Meter: 0141783,4 gallons

Toilet tank emptied: No

Deimos rover used: No, still not functional
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Not Applicable
Beginning charge: Not Applicable
Ending charge: Not Applicable
Currently charging: Not Applicable

Opportunity rover used: Used
Hours: 53,5 hours
Beginning charge: 100%
Ending charge: 74%
Currently charging: No

Curiosity rover used: Used
Hours: 81,8 hours
Beginning charge: 100%
Ending charge: 82%
Currently charging: Yes

Spirit rover used: Not Used
Hours: 82,7 hours
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Notes on rovers: Oppy is back on its wheels again ! No further issue with Curiosity to report.

ATV’s Used: N/A
Reason for use: Not used
Oil Added? N/A
ATV Fuel Used: N/A
# Hours the ATVs were Used today: None

Notes on ATVs: None

Hab Car used and why, where? Used by Staff.

Staff Crew Car used and why, where? Off Campus

General notes and comments: Nothing to report.

Summary of internet: Nothing to report.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report.

Summary of Science Dome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: See HSO Report.

Questions, concerns and requests to Mission Support: Nothing more to report.

Astronomy Report – Mar 15th

Astronomy Report, Sol 19

Name: Aurélien Mure

Crew: 206
Date: 03/15/2019

MUSK OBSERVATORY

Closing procedures done

Problem encountered : None

ROBOTIC OBSERVATORY

Object observed : Various Messier objects, including:

M100
M98
M104
M81
M66
M51
M101
M65
M82

Problems encountered : Clouds on most of the images, low exposure for
other ones, I will use MDRS WF and 14 tonight.

Astronomy Report – Mar 15th

Astronomy Report, Sol 19

Name: Aurélien Mure

Crew: 206
Date: 03/15/2019

MUSK OBSERVATORY

Closing procedures done

Problem encountered : None

ROBOTIC OBSERVATORY

Object observed : Various Messier objects

Problems encountered : Clouds on most of the images, low exposure for
other ones, I will use MDRS WF and 14 tonight.

Sol Summary Report – Mar 15th

Crew 206 March 15th 2019
Sol 19

Summary Title : Wrapping up the Sim

Author : Jeremy Auclair

Mission Status : All nominal

Sol Activity Summary : Today is our last “routine” day. We gave everything we had for the workout session, we ate quickly to be ready on time for the EVA which was earlier than usual. Half of the crew went exploring the ends of Galileo road while the other half started packing and continued cleaning the Hab, with only a few rest moments in the morning. The EVA teamed wrapped up the Weather station and LOAC disassembly around 10:30 am, and after they had arranged their equipment we started cleaning the lower deck. Everyone has a lot to do today, and we kept this fast pace all day long. At 11:30 am we started working on the daily reports and the preparation of lunch. After a copious meal, we went all around the station to pack our small equipment, film some last scenes and gather the last experimental
data. Even on this busy day we had to do the TELEOP session and the VR experiment, for the last time. The important end of mission reports are already done, so we are a little ahead of schedule. After the VR session, at 18h00, we opened the main airlock to breathe some fresh and absorb the sunlight. This is the end, we stayed outside for more than 45 minutes, taking in the surroundings for one last time. We are happy of the work we did, the experiences we lived and the memories we formed.

Look ahead plan : We are looking forward to meeting Crew 207, it will feel good to meet some new faces after 3 weeks of isolation. There is still work to do before leaving but the vast majority has been done. We will keep some personal time tomorrow morning to watch the sunrise from the Hab ridge, and perhaps jog towards the Mesa.

Anomalies in work : None

Weather : Sunny all day

Crew physical status : Fine

EVA : Weather station and LOAC disassembly and Galileo road exploration

Reports to be filed : GreenHab, Journalist, Operation, EVA Request, EVA Report

Support Requested : None